OpenStudy (anonymous):

Find the distance from the line y=2x+2 and the point 4,-6

OpenStudy (anonymous):

@phi @Michele_Laino @e.mccormick @nincompoop

OpenStudy (anonymous):

so i know you put 4,-6 into point slope form which is y+6=-2(x-4)

OpenStudy (anonymous):

then convert to slope incercept form which i think is y=-1/2x-4

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

NVM I GOT IT

OpenStudy (anonymous):

2x-y+2=0 distance of (x1,y1)from ax+by+c=0 is\[d=\left| \frac{ ax1+by1+c }{ \sqrt{a^2+b^2} } \right|\]